# StreamingTOM: Streaming Token Compression for Efficient Video Understanding

### v1

@@ Abstract
Unlike offline processing, streaming video vision-language models face two fundamental constraints: causality and accumulation. Causality prevents access to future frames that offline methods exploit, while accumulation causes tokens to grow unbounded, creating efficiency bottlenecks. However, existing approaches only regulate post-LLM kv-cache, leaving costly pre-LLM prefill unchanged. We introduce StreamingTOM, a training-free, plug-and-play two-stage framework that addresses both pre-LLM and post-LLM bottlenecks with predictable latency. Causal Temporal Reduction imposes a fixed per-frame budget and selects tokens based on adjacent-frame changes and token saliency, drastically reducing per-frame prefill cost by processing only a compact subset of visual tokens per frame instead of all visual tokens. Online Quantized Memory stores tokens in 4-bit format, retrieves relevant groups on demand, and dequantizes them, keeping the active kv-cache bounded regardless of stream length. Experiments demonstrate our method achieves $15.7\times$ kv-cache compression, $1.2\times$ lower peak memory and $2\times$ faster TTFT compared to prior SOTA. StreamingTOM maintains state-of-the-art accuracy among training-free methods with an average of $63.8\%$ on offline benchmarks and $55.8\%/3.7$ on RVS. These results highlight the practical benefits of our two-stage approach for efficient streaming video understanding with bounded growth.
@@
Figure 1: Left: StreamingTOM (streaming token compression) is a training-free, two-stage framework for efficient streaming video understanding: Causal Temporal Reduction selects pre-LLM tokens with a causal, fixed-budget policy, and Online Quantized Memory maintains compact kv-cache via 4-bit quantization with on-demand retrieval. Right: on LLaVA-OV-7B, StreamingTOM achieves a $15.7\times$ kv-cache compression ratio; compared to prior SOTA (LiveVLM), it delivers $1.2\times$ lower peak memory and $2\times$ faster TTFT.
@@ 1 Introduction
Recent advances in multimodal and large language models have demonstrated impressive capabilities across a diverse range of vision and language tasks [19, 11, 1, 36, 37]. Building on these foundations, offline video understanding investigates models for deep semantic reasoning on prerecorded videos with known temporal boundaries [2, 33, 3, 14, 45, 34]. A dominant efficiency bottleneck in these models is the computational cost of processing the heavy spatiotemporal redundancy in visual tokens. This motivated established compression techniques such as spatiotemporal token merging and dynamic pruning [25, 16], which leverage global context, including future frames, to reduce computation with little or no loss in accuracy [41, 29, 6, 24]. These techniques inherently assume access to a global context and stable clip boundaries.
While offline techniques have matured, emerging applications like autonomous driving, embodied AI, and live video assistants demand streaming processing, where two constraints emerge: causality and accumulation [15, 44, 35]. Causality means the system cannot access future frames. Accumulation arises because the effective sequence length $L(t)$ grows as frames arrive at a fixed frame rate. Dense attention grows quadratically with $L(t)$, while the kv-cache grows linearly with time [18]. As compute and memory grow unbounded over time, token compression shifts from an optional offline optimization to a fundamental prerequisite for streaming viability under fixed budgets.
@@
We propose StreamingTOM, a training-free framework that reduces visual tokens before the LLM for streaming video understanding. Our central insight is that temporal redundancy is abundant in both offline and streaming videos, but streaming combines it with strict causality and ever-growing context. To address this while respecting streaming constraints, effective compression must be pre-LLM and strictly causal. We impose a fixed per-frame budget $G$ to stabilize latency and ensure predictable compute and memory consumption across all frames. Accordingly, we implement this design with a strictly causal one-pass selector that makes decisions from adjacent frames t1t-1 and $t$. Causal Temporal Reduction (CTR) addresses the prefill bottleneck by selecting $G$ tokens per frame based on adjacent-frame changes and token saliency, reducing per-frame prefill cost by a factor of $N/G$ through processing only $G$ selected tokens instead of all $N$ visual tokens. [^1] Notation: NN=tokens per frame, LL=number of transformer layers, dd=hidden width, GG=per-frame retained tokens selected by CTR. Meanwhile, Online Quantized Memory (OQM) bounds post-LLM memory growth by storing retained tokens in 4-bit format and retrieving relevant groups on demand, keeping the active kv-cache bounded regardless of stream length. Together, these two stages enable long-horizon stability under fixed budgets with predictable latency, while remaining plug-and-play across different backbones.
StreamingTOM achieves $15.7\times$ kv-cache compression ratio and delivers $1.2\times$ lower peak memory and $2\times$ faster time to first token (TTFT) compared to prior training-free SOTA (LiveVLM [18]) on streaming setting. It maintains state-of-the-art accuracy among training-free methods with $63.8\%$ average on VideoMME [9], MLVU [46], and EgoSchema [17], and $55.8\%/3.7$ on RVS [44]. These gains persist over long horizons as our pre-LLM token reduction and bounded kv-cache prevent unbounded growth with video length. On LLaVA-OV-7B, a one-hour stream reduces kv-cache memory from $18.8\,\mathrm{GB}$ to $1.2\,\mathrm{GB}$, confirming bounded growth over extended sessions.
@@
  • We establish that effective streaming compression requires streaming-native design, introducing a strictly causal approach that reduces complexity from $O(tNLd^{2})$ to $O(tGLd^{2})$ with fixed per-frame budget $G$.
  • We propose StreamingTOM, a training-free two-stage framework combining CTR for pre-LLM compression with OQM for bounded post-LLM memory management.
  • We achieve state-of-the-art training-free accuracy ($63.8\%$ offline average, $55.8\%/3.7$ on RVS) with $15.7\times$ kv-cache compression, $1.2\times$ lower peak memory, and $2\times$ faster TTFT over training-free SOTA, demonstrating both effectiveness and efficiency.
@@ 2 Related Work
Training-free methods achieve substantial token reduction via hierarchical approaches [41, 38], adaptive pruning [12, 23], dynamic compression [26, 29], and holistic merging [24]. However, these approaches rely on global temporal analysis or bidirectional attention that leverages future frames. This look-ahead capability is incompatible with streaming causality, where compression must depend only on past context and cannot be revised. Training-based long-video compression methods also exist [27], but they require model retraining and are outside our training-free scope. Beyond token-level reduction, quantization techniques [30, 21, 31] offer memory compression through reduced precision, complementing spatial and temporal compression strategies.
@@
Training-based methods [4, 44, 40, 32, 22, 8, 39, 10, 5] achieve strong performance through response-timing control, memory decoupling, perception-decision disentanglement, multi-round interaction, visual feedback, and speech integration, but require costly model-specific retraining. Notably, TimeChat-Online [43] observes that 80% of visual tokens are redundant in streaming videos, motivating aggressive compression. In contrast, training-free approaches [7, 18, 42, 13] focus exclusively on post-LLM memory optimization through kv-cache retrieval, eviction, and compression strategies. However, all such methods preserve the full prefill computation cost because tokens have already traversed the full transformer stack, and the pre-LLM prefill remains intact.
To our knowledge, no prior training-free streaming method performs strictly causal pre-LLM token reduction. This gap is critical because pre-LLM compression cuts the prefill cost by reducing token count before the transformer, while post-LLM optimization only manages memory after the computational bottleneck. We address this with StreamingTOM, which couples Causal Temporal Reduction for pre-LLM compression with Online Quantized Memory for efficient kv-cache management under strict causality.
@@ 3.2 Our Method: StreamingTOM
Central to StreamingTOM is the group abstraction, a fixed-size unit of $G$ tokens per frame that serves as the interface between CTR and OQM, decoupling token reduction from storage optimization: CTR selects the most informative $G$ tokens, while OQM handles their efficient storage and retrieval. The fixed budget $G$ ensures predictable per-frame latency and enables stable batching across frames with varying content complexity. Overall, prefill computation reduces from $O(TNLd^{2})$ to $O(TGLd^{2})$, where $T$ denotes the total number of processed frames, and storage reduces from $O(TN\cdot d\cdot 16)$ bits to $O(TG\cdot d\cdot 4)$ bits, resulting in a combined compression ratio of $\frac{4N}{G}$.
@@ 3.3 Causal Temporal Reduction (CTR)
To partition tokens based on temporal similarity, CTR induces a measurable decomposition of the index space $[N]$ via threshold-parameterized indicators $I_{s}^{(i)},I_{d}^{(i)}:\mathbb{R}\to\{0,1\}$, yielding disjoint static/dynamic subsets,
$I_{s}^{(i)}=\mathbb{1}[s_{t}^{(i)}>\tau_{c}],\;I_{d}^{(i)}=1-I_{s}^{(i)},$ (5)
$\mathcal{S}_{t}=\{i\in[N]\mid I_{s}^{(i)}=1\},\;\mathcal{D}_{t}=[N]\setminus\mathcal{S}_{t}$
where $\tau_{c}$ is the similarity threshold. This deterministic partition enables compositional budget allocation. To compress from $N$ tokens to exactly $G$ tokens per frame, CTR allocates the budget proportionally based on the observed static/dynamic distribution,
@@
$\mathcal{G}_{d}=\text{TopK}_{k_{d}}(\{\alpha_{t}^{(i)}\}_{i\in\mathcal{D}_{t}}),\quad\mathcal{G}_{s}=\text{Merge}_{k_{s}}(\{\mathcal{F}_{t}^{(i)}\}_{i\in\mathcal{S}_{t}}).$ (7)
The per-frame retained group $\mathcal{G}_{t}$ combines both paths with $|\mathcal{G}_{t}|=G$. This fixed-size output is essential for streaming viability: unlike offline methods, where variable compression ratios are acceptable, streaming systems require predictable per-frame latency and stable batching to maintain real-time guarantees. CTR maintains only the previous-frame features as state, ensuring decisions are invariant to batch size and enabling consistent results across devices with different memory constraints. The state memory is $O(Nd)$ independent of stream length, preventing unbounded growth. Per-frame complexity is $O(N+G^{2})$ with $G<N$, dominated by the similarity computation. This results in a pre-LLM bound that reduces prefill from $O(TNLd^{2})$ to $O(TGLd^{2})$.
@@ 3.4 Online Quantized Memory (OQM)
where per-head, per-channel scale and offset are computed across tokens within each group, and a representative key $\bar{\mathbf{k}}_{t}\in\mathbb{R}^{d}$ is computed by averaging the keys before quantization to enable efficient retrieval,
$\boldsymbol{s}_{t,h,d}$ $=\frac{\max_{j\in[1,G]}\mathbf{X}_{t,h,j,d}-\min_{j\in[1,G]}\mathbf{X}_{t,h,j,d}}{15},$ (9)
$\boldsymbol{m}_{t,h,d}$ $=\min_{j\in[1,G]}\mathbf{X}_{t,h,j,d}$
and the 4-bit codes are packed into uint8 for storage. This yields a dual-structure memory: quantized groups with their representative keys are appended over time, while system tokens, such as instructions and task descriptions, remain in full precision,
$\mathcal{M}_{t}=\{\mathcal{KV}_{\mathrm{sys}}^{\mathrm{FP16}}\}\ \cup\ \{(\mathcal{Q}_{4}(\mathcal{G}_{i}),\bar{\mathbf{k}}_{i})\}_{i=1}^{t},$ (10)
and the per-frame storage ratio from FP16 to 4-bit can be computed as
@@ 4 Experiments
Table 1: Offline evaluation across three long-video benchmarks. Methods are grouped by design paradigm and training requirement, with “+” indicating LLaVA-OV-7B implementations. The best and second best results within the streaming-designed training-free category are in bold and underlined, respectively.
Method Frames VideoMME MLVU EgoSchema Avg.
Long Medium Short Overall
@@ 4.1 Experimental Setups
Pre-LLM reduction with CTR retains 50 tokens per frame based on adjacent-frame changes and token saliency, using a strictly causal two-frame decision without look-ahead; we set the similarity threshold to 0.9 to determine frame-level redundancy. Post-LLM memory with OQM stores retained tokens in 4-bit quantized groups and retrieves relevant groups on demand with a total budget of 12k tokens. For fair comparison with ReKV, we align this budget to match ReKV baseline’s $64$ frames at $196$ tokens per frame. Cached tokens are organized into groups of size 50 to stabilize memory. The streaming encoder batch size defaults to 32; our compression is batch-agnostic and supports flexible deployment.
@@ 4.2 Main Results
In Table 1, StreamingTOM achieves the best overall performance among all streaming systems, regardless of whether they require training. Our training-free method reaches an average of 63.8, surpassing the strongest training-free baseline StreamMem [42] at 63.1 and outperforming training-based Dispider-7B. StreamingTOM demonstrates consistently strong performance across individual benchmarks, achieving 59.9 on VideoMME-Overall, 67.9 on MLVU, and 63.7 on EgoSchema, consistently outperforming the training-free baseline StreamMem. Within VideoMME splits, StreamingTOM excels particularly on Long and Medium subsets while showing comparable performance on Short videos, demonstrating robust capability across different temporal scales. These results clearly show that under the same backbone and protocol, StreamingTOM sets a new SOTA for streaming video understanding despite being training-free.
@@
We evaluate StreamingTOM on RVS-Ego and RVS-Movie, using GPT-3.5-turbo-0125 [20] for grading and limiting GPU memory to 28  GB without CPU offloading for fair comparison. Under this memory-constrained setting, StreamingTOM achieves the strongest overall performance with 55.8 Acc and 3.7 Score, surpassing the strongest training-free baseline StreamMem at 55.2/3.6 and outperforming other methods including InfiniPot-V, ReKV, and training-based Flash-VStream. While ReKV with CPU offloading reports higher accuracy (59.0), this strategy relies on CPU-GPU memory transfers that are impractical for real-time streaming deployment. We therefore compare under memory-constrained settings without CPU offloading, where StreamingTOM achieves the strongest performance. StreamingTOM achieves 58.3/3.9 on RVS-Ego and 53.2/3.5 on RVS-Movie, demonstrating consistent performance across both. We further provide qualitative examples on RVS Movie (Figure 4), where StreamingTOM produces faithful answers aligned with ground truth, demonstrating its ability to capture fine-grained semantics under causal streaming constraints.

### v2

@@ Abstract
Unlike offline processing, streaming video vision-language models face two fundamental constraints: causality and accumulation. Causality prevents access to future frames that offline methods exploit, while accumulation causes tokens to grow unbounded, creating efficiency bottlenecks. However, existing approaches only regulate post-LLM kv-cache, leaving costly pre-LLM prefill unchanged. We introduce StreamingTOM, a training-free, plug-and-play two-stage framework that addresses both pre-LLM and post-LLM bottlenecks. Causal Temporal Reduction imposes a fixed per-frame budget and selects tokens based on adjacent-frame changes and token saliency, drastically reducing per-frame prefill cost by processing only a compact subset of visual tokens, ensuring predictable latency. Online Quantized Memory stores tokens in 4-bit format, retrieves relevant groups on demand, and dequantizes them, keeping the active kv-cache bounded regardless of stream length. Experiments demonstrate our method achieves $15.7\times$ kv-cache compression ratio; compared to prior SOTA (LiveVLM), it delivers $1.2\times$ lower peak memory and $2\times$ faster TTFT. StreamingTOM achieves state-of-the-art accuracy among training-free methods with an average of $63.8\%$ on offline benchmarks and $55.8\%$ accuracy and $3.7$ score on RVS. These results demonstrate that real-time streaming video understanding with bounded active memory is achievable without model retraining.
@@
Figure 1: Left: StreamingTOM (streaming token compression) is a training-free, two-stage framework for efficient streaming video understanding: Causal Temporal Reduction selects pre-LLM tokens with a causal, fixed-budget policy, and Online Quantized Memory bounds the active kv-cache via 4-bit quantization with on-demand retrieval. Right: On LLaVA-OV-7B, StreamingTOM achieves a $15.7\times$ kv-cache compression ratio; compared to prior SOTA (LiveVLM), it delivers $1.2\times$ lower peak memory and $2\times$ faster TTFT.
@@ 1 Introduction
Recent advances in large language models have significantly improved both reasoning capabilities [10, 35, 22, 39, 63, 16, 29, 58, 59, 13] and multimodal understanding [34, 17, 1, 19], demonstrating impressive performance across diverse vision and language tasks. Building on these foundations, offline video understanding focuses on deep semantic reasoning over prerecorded videos with known temporal boundaries [2, 52, 3, 23, 71, 53]. A dominant efficiency bottleneck in these models is the computational cost of processing the spatiotemporal redundancy in visual tokens [20]. This motivated established compression techniques such as spatiotemporal token merging and dynamic pruning [43, 25], which leverage global context, including future frames, to reduce computation with little or no loss in accuracy [65, 47, 8, 42]. These techniques inherently assume access to a global context and stable clip boundaries.
While offline techniques have matured, emerging applications like autonomous driving, embodied AI, and live video assistants demand streaming processing, where two constraints emerge: causality and accumulation [24, 69, 56]. Causality means the system cannot access future frames. Accumulation arises because the effective sequence length $L(t)$ grows as frames arrive at a fixed frame rate. Dense attention grows quadratically with $L(t)$, while the kv-cache grows linearly with time [32]. As compute and memory grow unbounded over time, token compression shifts from an optional optimization to a fundamental prerequisite for streaming viability under fixed budgets.
@@
We propose StreamingTOM, a training-free framework that reduces visual tokens before the LLM for streaming video understanding. Our central insight is that effective streaming compression must act before the LLM under strict causality: post-LLM methods cannot reduce the already-incurred prefill cost, while offline pre-LLM methods violate causality by leveraging future frames. We impose a fixed per-frame budget $G$ to stabilize latency and ensure predictable compute and memory consumption across all frames. Accordingly, we implement this design with a strictly causal one-pass selector that makes decisions from adjacent frames t1t-1 and $t$. Causal Temporal Reduction (CTR) addresses the prefill bottleneck by selecting $G$ tokens per frame based on adjacent-frame changes and token saliency, reducing per-frame prefill cost by a factor of $N/G$ through processing only $G$ tokens instead of all $N$ visual tokens. [^1] Notation: NN=tokens per frame, LL=number of transformer layers, dd=hidden width, GG=per-frame retained tokens selected by CTR. Meanwhile, Online Quantized Memory (OQM) bounds post-LLM memory growth by storing retained tokens in 4-bit format and retrieving relevant groups on demand, keeping the active kv-cache bounded regardless of stream length. Together, these two stages enable long-horizon stability under fixed budgets with predictable latency, while remaining plug-and-play across backbones.
StreamingTOM achieves $15.7\times$ kv-cache compression ratio and delivers $1.2\times$ lower peak memory and $2\times$ faster time to first token (TTFT) compared to prior training-free SOTA (LiveVLM [32]) in the streaming setting. It maintains state-of-the-art accuracy among training-free methods with $63.8\%$ average on VideoMME [14], MLVU [72], and EgoSchema [31], and $55.8\%/3.7$ on RVS [69]. These gains persist over long horizons because CTR’s fixed budget caps per-frame cost while OQM bounds the active kv-cache independently of stream length. On LLaVA-OV-7B, a one-hour stream reduces kv-cache from $18.8\,\mathrm{GB}$ to $1.2\,\mathrm{GB}$, confirming bounded growth over extended sessions.
@@
  • We introduce strictly causal pre-LLM token reduction for streaming video understanding, reducing prefill complexity from $O(tNLd^{2})$ to $O(tGLd^{2})$ through a fixed per-frame budget $G$ that ensures predictable latency.
  • We propose StreamingTOM, a training-free two-stage framework unified by a frame-aligned group abstraction of $G$ tokens that coordinates CTR for pre-LLM compression with OQM for bounded post-LLM memory, ensuring temporal coherence across compression and retrieval.
  • We achieve state-of-the-art training-free accuracy ($63.8\%$ offline average, $55.8\%/3.7$ on RVS) with $15.7\times$ kv-cache compression ratio, $1.2\times$ lower peak memory, and $2\times$ faster TTFT over training-free SOTA, demonstrating both effectiveness and efficiency.
@@ 2 Related Work
Training-free methods achieve substantial token reduction via spatial selection and adaptive pruning [4, 40, 65, 60, 18, 41, 6, 64], and dynamic spatio-temporal compression and merging [44, 47, 48, 26, 28, 42, 55]. However, these approaches rely on global temporal analysis or bidirectional attention that leverages future frames. This look-ahead capability is incompatible with streaming causality, where compression must depend only on past context and cannot be revised. Training-based long-video compression methods also exist [45], but they require model retraining and are outside our training-free scope. Beyond token-level reduction, complementary compression strategies such as model pruning [30, 74], kv-cache quantization [49, 36, 50, 27], and tokenized video compression [73] further reduce memory footprint from orthogonal dimensions.
@@
Training-based methods [5, 69, 62, 51, 37, 38, 12, 61, 15, 7] achieve strong performance through response-timing control, memory decoupling, and perception-decision disentanglement, but require costly model-specific retraining. TimeChat-Online [67] observes that 80% of visual tokens are redundant in streaming videos, motivating aggressive compression. In contrast, training-free approaches [11, 32, 66, 21] focus on post-LLM memory optimization through kv-cache retrieval, eviction, and compression strategies, leaving the full prefill computation cost intact. Concurrent work [54] also explores token compression for streaming video LLMs.
To our knowledge, no prior training-free streaming method performs strictly causal pre-LLM token reduction. We address this with StreamingTOM, which couples CTR for pre-LLM compression with OQM for kv-cache management under strict causality.
@@ 3.2 Our Method: StreamingTOM
Central to StreamingTOM is the group abstraction, a fixed-size unit of $G$ tokens per frame that serves as the interface between CTR and OQM, decoupling token reduction from storage optimization: CTR selects the most informative $G$ tokens, while OQM handles their efficient storage and retrieval. This frame-aligned design is critical for streaming: each group corresponds to exactly one frame, ensuring that retrieved content preserves temporal coherence as complete frames rather than fragmented tokens. Without this coordination, variable-length compression would break frame-aligned retrieval, and token-level retrieval would lose temporal semantics. The fixed budget $G$ further ensures predictable per-frame latency and enables stable batching across frames with varying content complexity. Overall, prefill computation reduces from $O(TNLd^{2})$ to $O(TGLd^{2})$, where $T$ denotes the total number of processed frames, and storage reduces from $O(TN\cdot d\cdot 16)$ bits to $O(TG\cdot d\cdot 4)$ bits, resulting in a combined compression ratio of $\frac{4N}{G}$. [^3] With $N=196$ and $G=50$, this gives $4\times 196/50\approx 15.7\times$.
@@ 3.3 Causal Temporal Reduction (CTR)
CTR partitions tokens into disjoint static and dynamic subsets by thresholding the temporal similarity,
$\mathcal{S}_{t}=\{i\in[N]\mid s_{t}^{(i)}>\tau_{c}\},\qquad\mathcal{D}_{t}=[N]\setminus\mathcal{S}_{t},$ (5)
where $\tau_{c}$ is the similarity threshold. To compress from $N$ tokens to exactly $G$ tokens per frame, CTR allocates the budget proportionally based on the observed static/dynamic distribution,
@@
$\mathcal{G}_{d}$ $=\text{TopK}_{k_{d}}(\{\alpha_{t}^{(i)}\}_{i\in\mathcal{D}_{t}}),$ (7)
$\mathcal{G}_{s}$ $=\text{Merge}_{k_{s}}(\{\mathcal{F}_{t}^{(i)}\}_{i\in\mathcal{S}_{t}}).$
The per-frame retained group $\mathcal{G}_{t}$ combines both paths with $|\mathcal{G}_{t}|=G$. This fixed-size output is essential for streaming viability: unlike offline methods, where variable compression ratios are acceptable, streaming systems require predictable per-frame latency and stable batching to maintain real-time guarantees. CTR maintains only the previous-frame features as state, ensuring decisions are invariant to batch size and enabling consistent results across devices with different memory constraints. The state memory is $O(Nd)$ independent of stream length, preventing unbounded growth. Per-frame complexity is $O(N+G^{2})$ with $G<N$, dominated by the clustering computation. This results in a pre-LLM bound that reduces prefill from $O(TNLd^{2})$ to $O(TGLd^{2})$.
@@ 3.4 Online Quantized Memory (OQM)
where the scale and offset are computed per-head, per-channel across the $G$ tokens in each group,
$\boldsymbol{s}_{t}=\frac{\max(\mathbf{X}_{t})-\min(\mathbf{X}_{t})}{15},\qquad\boldsymbol{m}_{t}=\min(\mathbf{X}_{t}),$ (9)
and a representative key $\bar{\mathbf{k}}_{t}\in\mathbb{R}^{d}$ is obtained by averaging the keys before quantization. The 4-bit codes are packed into uint8 for storage. This yields a dual-structure memory $\mathcal{M}_{t}$: system tokens (instructions and task descriptions) remain in FP16, while quantized visual groups $\mathcal{Q}_{4}(\mathcal{G}_{i})$ with their representative keys $\bar{\mathbf{k}}_{i}$ are appended incrementally over time. The per-frame storage ratio from FP16 to 4-bit can be computed as
@@ 4 Experiments
Table 1: Offline evaluation across three long-video benchmarks. “+” denotes LLaVA-OV-7B implementations. Best and second best in the streaming training-free category.
Method Fr. VideoMME MLVU Ego. Avg.
L M S Ovl.
@@ 4.1 Experimental Setups
Pre-LLM reduction with CTR retains 50 tokens per frame based on adjacent-frame changes and token saliency, using a strictly causal two-frame decision without look-ahead. We set the similarity threshold to 0.9 to determine frame-level redundancy. Post-LLM memory with OQM stores retained tokens in 4-bit quantized groups and retrieves relevant groups on demand with a total budget of 12k tokens. For fair comparison with ReKV, we align this budget to match ReKV baseline’s $64$ frames at $196$ tokens per frame. Cached tokens are organized into groups of size 50 to stabilize memory. The streaming encoder batch size defaults to 32; our compression is batch-agnostic and supports flexible deployment.
@@ 4.2 Main Results
In Table 1, StreamingTOM achieves the best overall performance among all streaming systems, regardless of whether they require training. Our training-free method reaches an average of 63.8, surpassing the strongest training-free baseline StreamMem [66] at 63.1 and outperforming training-based Dispider-7B. StreamingTOM demonstrates strong performance across individual benchmarks, achieving 59.9 on VideoMME-Overall, 67.9 on MLVU, and 63.7 on EgoSchema. Within VideoMME splits, StreamingTOM excels particularly on Long and Medium subsets while showing comparable performance on Short videos, demonstrating robust capability across different temporal scales. These results clearly show that under the same backbone and evaluation protocol, StreamingTOM sets a new SOTA for streaming video understanding despite being entirely training-free.
@@
We evaluate StreamingTOM on RVS-Ego and RVS-Movie, using GPT-3.5-turbo-0125 [33] for grading and limiting GPU memory to 28  GB without CPU offloading for fair comparison. Under this memory-constrained setting, StreamingTOM achieves the strongest overall performance with 55.8 Acc and 3.7 Score, surpassing the strongest training-free baseline StreamMem at 55.2/3.6 and outperforming other methods including InfiniPot-V, ReKV, and training-based Flash-VStream. While ReKV with CPU offloading reports higher accuracy (59.0), this strategy relies on CPU-GPU memory transfers that are impractical for real-time streaming deployment. We therefore compare under memory-constrained settings without CPU offloading. StreamingTOM achieves 58.3/3.9 on RVS-Ego and 53.2/3.5 on RVS-Movie, demonstrating consistent performance across both. We further provide qualitative examples on RVS Movie (Figure 4), where StreamingTOM produces faithful answers aligned with ground truth, demonstrating its ability to capture fine-grained semantics under causal streaming constraints.